# Higher-Dimensional Moving Averages and Submanifold Genericity

### v1

@@ 1.4. Acknowledgments
As in the proof of 1.5, take $\alpha=1$ and suppose that ($!\tilde{C}_{i}$) holds for $i=1$. Fix $p\in\mathbb{N}$ and choose a real number $\lambda_{p}>0$ so that $\mathrm{Leb}\left(\Omega_{1}^{(1)}(\lambda_{p})\right)\geq p\cdot 4\lambda_{p}$. Then for each $z\in\Omega_{1}^{(1)}(\lambda_{p})$ we have that
@@
$Q_{p}:=[L_{1}-4\lambda_{p},L_{1})\times[0,3L_{2})\times\dotsm\times[0,3L_{d})\vspace{2mm}$
@@
$\frac{\mu\left(N_{\mathcal{B}}\chi_{H_{p}}>1-\varepsilon\right)}{\mu(H_{p})}\geq\frac{p\cdot 4\lambda_{p}}{2^{d-1}\cdot L_{1}}\cdot\frac{L_{1}}{4\lambda_{p}}=\frac{p}{2^{d-1}},$
showing ($\star$3.1) for the operator ${N_{\mathcal{B}}}$. To conclude, note that if :=$\mathcal{B}^{\prime}:=\{B_{k}\in\mathcal{B}:k\geq n\}$, then ($!\tilde{C}_{i}$) holds for the collection $\mathcal{B}^{\prime}$, and the above argument also applies to the operator $N_{\mathcal{B}^{\prime}}$.
@@
Now, let us consider a new action of $\mathbb{R}^{m+1}$ on $X$ (which we denote by ".."".."" . . "), defined by the relations:
@@
Then we have that
@@
Note that, in the notation of Theorem 1.9 we have that $s_{k1}=1$ for all $k$. Thus, by Remark 2.3, Condition ($!\tilde{C}_{i}$) holds for $i=1$. Moreover, the action of $\mathbb{R}^{m+1}$ defined above is aperiodic, since $\boldsymbol{u},\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{m}$ are linearly independent, and it commutes with a mixing family of transformations on $X$, by Remark 1.6. Thus, by Theorem 1.9, for any $\varepsilon>0$ there exist a measurable set $E_{\varepsilon}\subset X$ and a sequence of integers $k_{r}$ such that for all $r$ and $\mu$-a.e. $x\in X$ it holds that
$\frac{1}{k_{r}^{m}}\int_{[k_{r}-1,k_{r})\times[0,k_{r})^{m}}\chi_{E_{\varepsilon}}\Big{(}(\lambda_{0}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big{)}\,\textup{d}\lambda_{0}\textup{d}\boldsymbol{\lambda}>97/100.$
@@
$\frac{1}{k_{r}^{m}}\int_{[0,k_{r})^{m}}\chi_{E_{\varepsilon}}\Big{(}(t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big{)}\textup{d}\boldsymbol{\lambda}>\frac{97}{100},$
@@
$\frac{1}{k_{r}^{m}}\int_{(0,t_{r})^{m}}\chi_{E_{\varepsilon}}\Big{(}(t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big{)}\textup{d}\boldsymbol{\lambda}\\ =\frac{1}{k_{r}^{m}}\int_{[0,k_{r})^{m}}\chi_{E_{\varepsilon}}\Big{(}(t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big{)}\textup{d}\boldsymbol{\lambda}+O_{m}(k_{r}^{-1})>\frac{97}{100}+O_{m}(k_{r}^{-1}).$
@@
(4.5) $\frac{1}{t_{r}^{m}}\int_{(0,t_{r})^{m}}\chi_{E_{\varepsilon}}\Big{(}(t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb\lambda_{m}\boldsymbol{e}_{m})..x\Big{)}\textup{d}\boldsymbol{\lambda}\\ >\frac{97k_{r}^{m}}{100t_{r}^{m}}+O_{m}(t_{r}^{-1})=\frac{97}{100}+O_{m}(t_{r}^{-1}).$
@@
$\frac{1}{t_{r}^{m}\cdot\textup{vol}_{m}(M)}\int_{t_{r}M}\chi_{E}(\boldsymbol{a}.x)\,\textup{d}\textup{vol}_{m}(\boldsymbol{a})\geq\frac{97\sqrt{\det\left(V^{T}V\right)}}{100\textup{vol}_{m}(M)}+O_{m,M,V}(t_{r}^{-1}).$

### v2

@@ 1.3. Acknowledgments
Take $\alpha=1$ and suppose that ($\tilde{C}_{i}$) holds for $i=1$. Fix $p\in\mathbb{N}$ and choose a real number $\lambda_{p}>0$ so that $\mathrm{Leb}\left(\Omega_{1}^{(1)}(\lambda_{p})\right)\geq p\cdot 4\lambda_{p}$. Then for each $z\in\Omega_{1}^{(1)}(\lambda_{p})$ we have that
@@
$Q_{p}:=[L_{1}-4\lambda_{p},L_{1})\times[0,3L_{2})\times\dotsm\times[0,3L_{d})\vskip 5.69054pt$
@@
$\frac{\mu\left(N_{\mathcal{B}}\chi_{H_{p}}>1-\varepsilon\right)}{\mu(H_{p})}\geq\frac{p\cdot 4\lambda_{p}}{3^{d-1}\cdot L_{1}}\cdot\frac{L_{1}}{4\lambda_{p}}=\frac{p}{3^{d-1}},$
showing ($\star$3.1) for the operator $N_{\mathcal{B}}$}. To conclude, note that if $\mathcal{B}^{\prime}:=\{B_{k}\in\mathcal{B}:k\geq n\}$, then ($\tilde{C}_{i}$) holds for the collection $\mathcal{B}^{\prime}$, and the above argument also applies to the operator $N_{\mathcal{B}^{\prime}}$.
@@
Now, let us consider a new action of $\mathbb{R}^{m+1}$ on $X$ (which we denote by ..``.."), defined by the relations:
@@
Then for $\boldsymbol{\lambda}=(\lambda_{1},\dotsc,\lambda_{m})$ we have that
@@
Note that, in the notation of Theorem 1.5, we have that $s_{k1}=1$ for all $k$. Thus, by Remark 2.3, Condition ($\tilde{C}_{i}$) holds for $i=1$. Moreover, the action of $\mathbb{R}^{m+1}$ defined above is aperiodic, since $\boldsymbol{u},\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{m}$ are linearly independent, and it commutes with a mixing family of transformations on $X$, by Remark 1.6. Thus, by Theorem 1.5, for any $\varepsilon>0$ there exist a measurable set $E_{\varepsilon}\subset X$ and a sequence of integers $k_{r}$ such that for all $r$ and $\mu$-a.e. $x\in X$ it holds that
$\frac{1}{k_{r}^{m}}\int_{[k_{r}-1,k_{r})\times[0,k_{r})^{m}}\chi_{E_{\varepsilon}}\Big((\lambda_{0}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big)\,\textup{d}\lambda_{0}\textup{d}\boldsymbol{\lambda}>0.99.$
@@
$\frac{1}{k_{r}^{m}}\int_{[0,k_{r})^{m}}\chi_{E_{\varepsilon}}\Big((t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big)\textup{d}\boldsymbol{\lambda}>0.99,$
@@
$\frac{1}{k_{r}^{m}}\int_{(0,t_{r})^{m}}\chi_{E_{\varepsilon}}\Big((t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big)\textup{d}\boldsymbol{\lambda}\\ =\frac{1}{k_{r}^{m}}\int_{[0,k_{r})^{m}}\chi_{E_{\varepsilon}}\Big((t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb+\lambda_{m}\boldsymbol{e}_{m})..x\Big)\textup{d}\boldsymbol{\lambda}+O_{m}(k_{r}^{-1})>0.99+O_{m}(k_{r}^{-1}).$
@@
(4.5) $\frac{1}{t_{r}^{m}}\int_{(0,t_{r})^{m}}\chi_{E_{\varepsilon}}\Big((t_{r}\boldsymbol{e}_{0}+\lambda_{1}\boldsymbol{e}_{1}+\dotsb\lambda_{m}\boldsymbol{e}_{m})..x\Big)\textup{d}\boldsymbol{\lambda}\\ >0.99\cdot\frac{k_{r}^{m}}{t_{r}^{m}}+O_{m}(t_{r}^{-1})=0.99+O_{m}(t_{r}^{-1}).$
@@
$\frac{1}{t_{r}^{m}\cdot\textup{vol}_{m}(M)}\int_{t_{r}M}\chi_{E}(\boldsymbol{a}.x)\,\textup{d}\textup{vol}_{m}(\boldsymbol{a})\geq 0.99\cdot\frac{\sqrt{\det\left(V^{T}V\right)}}{\textup{vol}_{m}(M)}+O_{m,M,V}(t_{r}^{-1}).$